# VibeCodeHPC: An Agent-Based Iterative Prompting Auto-Tuner for HPC Code Generation Using LLMs

### v2

@@ I Introduction
In recent years, large language models (LLMs) have demonstrated remarkable capabilities in code generation, automatically transforming natural language specifications into executable programs. These models, trained on large-scale source code corpora and natural language texts, have been applied to various aspects of software development, including rapid prototyping, code refactoring, and automated testing. More recently, coding agent systems have emerged that autonomously accomplish tasks initially specified by users through automatic iterative operations. Representative systems include Claude Code [^1] https://github.com/anthropics/claude-code, Codex CLI [^2] https://github.com/openai/codex, and Gemini CLI [^3] https://github.com/google-gemini/gemini-cli. These systems have become even more user-friendly tools for developers through integration with command-line interfaces (CLI). However, while LLM-based systems have shown promising results in generating functionally correct code, their application to performance-critical domains such as high-performance computing (HPC) applications executed on supercomputers remains limited.
HPC programming involves stringent requirements beyond functional correctness. For example, it requires the application of architecture-specific optimization techniques, achieving high performance particularly through thread and process parallelization, minimizing power consumption, and balancing numerical computation precision with computational speed. Additionally, HPC software typically utilizes parallel programming models such as MPI (Message Passing Interface), OpenMP, CUDA, and OpenACC. These models must be appropriately selected based on application characteristics and the target execution environment. Furthermore, utilizing supercomputer systems requires system-specific operations that differ significantly from program development on typical PCs. Specifically, this includes operating batch job systems, managing software modules, and managing Secure Shell (SSH) connections.
The complex development tasks described above pose difficult challenges not only for humans but also for code generation AI agents. Similar to humans, LLMs have memory capacity constraints (context windows), and when a single agent performs complex tasks for extended periods, situations may arise where it forgets initially given instructions. In such circumstances, the effectiveness of multi-agent systems where multiple LLMs with specialized roles operate collaboratively is expected. By having each agent focus on a limited role, it may be possible to reduce the memory burden on individual agents and reliably accomplish more complex tasks. This effect is considered particularly pronounced in highly complex development such as HPC programming. Moreover, in complex development such as HPC programming, creating rigorous requirement definitions in advance is also a difficult task for humans. Recently, a development paradigm called Vibe Coding [19][28] has been proposed, where programs are incrementally developed through dialogue between humans and LLMs without relying on strict requirement definitions. The term “Vibe” in Vibe Coding carries connotations of “getting into the groove” and “feeling the atmosphere,” with dialogue between humans and LLMs occurring as if it were informal communication. Through this dialogue, programs are gradually shaped. Such a development paradigm may be particularly effective for complex development targets where creating rigorous requirement definitions is difficult, such as HPC programming.
@@
Based on the above background, this study proposes VibeCodeHPC, a multi-agent system based on LLMs for the automatic tuning of HPC programs on supercomputers. VibeCodeHPC adopts Claude Code as its backend and provides an integrated environment that facilitates program development in supercomputer settings. Thanks to its sophisticated multi-agent mechanism, the system not only enables vibe coding as if working with a “specialized HPC programming team,” but also achieves fully automated tuning through autonomous operation that minimizes user intervention. To realize these objectives, VibeCodeHPC provides the following core functionalities:
@@
The structure of this paper is as follows. Section II reviews related work. Section III presents an overview of VibeCodeHPC and its implementation, with a particular focus on its sophisticated multi-agent features. Section IV demonstrates one of the most powerful features and challenges of VibeCodeHPC: fully automated code optimization through autonomous operation without user intervention. Specifically, it evaluates the performance optimization of CPU-based codes on GPU-equipped systems for matrix multiplication and a Poisson equation solver using Jacobi’s iterative method. Finally, Section V summarizes the findings of this study and discusses future research directions.
@@ II Related Work
II-A LLM-Based Code Generation
Large language models (LLMs) including GPT-4, LLaMA 3, and Code LLaMA have significantly advanced automatic code generation. These models have demonstrated capabilities in producing syntactically correct and semantically meaningful code from natural language descriptions, with applications in software prototyping, automated bug fixing, and code refactoring. Chen et al. [4] showed that models trained on large-scale code corpora can generate effective solutions for programming contest problems. Austin et al. [2] and Li et al. [26] introduced prompt engineering and in-context learning strategies to improve code correctness and robustness. However, these studies primarily focus on general-purpose programming tasks without addressing HPC-specific performance constraints.
II-B HPC Code Generation and Optimization
@@
HPC code generation traditionally relies on domain experts to manually implement performance-critical kernels using parallel programming models such as MPI, OpenMP, CUDA, and OpenACC. Compiler-assisted approaches [43] [23] have aimed to automate parts of this process by applying loop transformations, memory layout optimizations, and vectorization. More recent studies have investigated domain-specific languages (DSLs) [40] [16] to generate optimized HPC kernels from high-level specifications. While these methods can yield high performance, they typically require explicit user annotations or DSL expertise, limiting their accessibility. The integration of LLMs into HPC code generation remains largely unexplored, especially in conjunction with automated performance tuning.
@@
LLM-based HPC code generation. Recent studies have begun evaluating LLMs for parallel programming models and HPC kernels, including OpenMP-centric models and MPI+OpenMP code generation in scalable workloads [3] [7]. Early evaluations of LLMs for OpenMP/MPI/CUDA/OpenACC kernels [10] [38] and LLM-assisted portability from Fortran to C/C++ [37] suggest that HPC-specific constraints substantially affect correctness and performance outcomes. These works provide concrete baselines against which VibeCodeHPC can be positioned.
@@
HPC operations and applied studies. Practical case studies that combine LLM-based coding agents with GPU acceleration and testing pipelines offer direct evidence of feasibility in HPC contexts [14]. Additionally, HPC-oriented LLM systems and evaluations for software development provide a broader landscape for integration and operational support [46, 8, 11, 30].
II-C AT and Iterative Refinement
@@
AT has been widely adopted in HPC to search parameter spaces and identify optimal configurations for target architectures. Frameworks such as ATLAS [42], OpenTuner [1], and GPTune [27] automate this process, often leveraging search algorithms like Bayesian optimization, genetic algorithms, or multi-armed bandits. Additionally, iterative refinement strategies—where code is repeatedly modified based on performance feedback—have been explored in machine learning-guided compilation [5] [12] . While these approaches are effective in optimizing existing code, their integration with LLM-driven code synthesis is still in its infancy.
Agentic optimization vs autotuning. Comparative analyses that pit agentic LLM approaches against classical autotuning for loop transformations suggest that LLM-driven strategies can complement or accelerate parameter search in certain settings [34]. This line of work motivates combining structured tuning with multi-agent orchestration.
II-D Iterative Prompt in LLM-based Code Generation
@@
However, iteration can introduce vulnerabilities. Shukla et al. [35] reported a 37.6% increase in critical vulnerabilities after five iterations, highlighting the need for human oversight. From a structural perspective, Ren et al. [32] proposed Knowledge-driven Prompt Chaining (KPC) for exception handling, while Li et al. [25] introduced Structured CoT, improving Pass@1 accuracy by up to 13.79% on HumanEval. Ridnik et al. [33] demonstrated that test-driven, multi-stage iterative flows improved Pass@5 accuracy from 19% to 44%.
Wei et al. [41] investigated CUDA code optimization using multi-agent systems without additional training. However, the role-sharing approach proposed in our study has not been evaluated for cross-language optimization.
These studies constitute methodological investigations into iterative prompting. Research on AI-agent collaboration paradigms, as exemplified by VibeCodeHPC, remains nascent.
Tech trends. Beyond research papers, the recent surge of general-purpose multi-agent frameworks (e.g., AutoGen[44], MetaGPT[13], CAMEL[24], and Magentic-One[9]) and agent orchestration tools (Claude Code subagents [^4] https://docs.anthropic.com/en/docs/claude-code/sub-agents, SuperClaude [^5] https://github.com/NomenAK/SuperClaude, Claude-Flow [^6] https://github.com/ruvnet/claude-flow, Claude Squad [^7] https://github.com/smtg-ai/claude-squad, Claude Task Master [^8] https://github.com/eyaltoledano/claude-task-master, Vibe Kanban [^9] https://github.com/BloopAI/vibe-kanban, KAMUI [^10] https://3rd.kamui.ai/en) highlights the growing interest in collaborative LLM workflows. Microsoft’s agent orchestration patterns categorize Magentic orchestration as a distinct mode, which aligns closely with VibeCodeHPC’s collaborative, role-separated workflow; however, a framework explicitly specialized for supercomputer code development and operations remains unproposed. [29].
@@ II-F Summary
Existing research in LLM-based code generation has achieved significant progress in producing functionally correct programs, but it rarely addresses HPC-specific performance optimization. On the other hand, the HPC community has developed powerful auto-tuning and compiler-based optimization techniques, yet these methods often lack the flexibility and adaptability of LLM-based code synthesis. Bridging these domains—by integrating LLM-driven generation, HPC performance tuning, and iterative improvement—remains an open challenge, which our proposed VibeCodeHPC framework seeks to address.
@@ III Overview and Implementation of VibeCodeHPC
This section explains the system overview of VibeCodeHPC, the operational flow from the user’s perspective, and the implementation that realizes it. In particular, we provide detailed explanation of the multi-agent implementation, which is the core innovation of VibeCodeHPC.
@@ III-A System Components
VibeCodeHPC consists of the following three main components.
@@
    • SSH connection management: Establishment and maintenance of remote access to supercomputers (Figure 1)
@@
    This information is provided to Claude Code to support agents in appropriately performing system-specific operations.
@@
Figure 1: SSH connection architecture. Private keys are managed by ssh-agent and not exposed to the LLM context. Desktop Commander MCP server maintains persistent SSH/SFTP sessions.
@@ III-B Workflow
The typical usage flow of VibeCodeHPC is shown below.
@@
    System Startup: VibeCodeHPC can be operated anywhere, but considering the typical usage pattern of supercomputers, it is assumed that VibeCodeHPC operates on the user’s local terminal (Linux environment) that accesses the supercomputer. The user starts VibeCodeHPC on the local terminal, and VibeCodeHPC launches Claude Code.
@@
    Monitoring and Result Confirmation: Users can monitor the operation of each agent in real time. Additionally, VibeCodeHPC automatically generates work logs, allowing users to grasp optimization progress by reviewing them. When VibeCodeHPC achieves the user’s objectives, a final report is automatically generated. This report includes the optimization process, achieved performance, techniques used, etc.
@@ III-C1 Agent Configuration
VibeCodeHPC enables collaborative operation of multiple agents specialized for different tasks. Specifically, the following agents are defined. The roles of these agents are defined by markdown-format configuration files that Claude Code loads at startup, and users can customize them.
@@
  • Programmer (PG): An agent responsible for actual programming tasks. PGs are deployed for each specialized field. Examples of specialized fields include CPU optimization, GPU optimization (CUDA, OpenACC), distributed parallelization using MPI, and algorithm optimization. This deployment has templates prepared according to the system configuration of the target supercomputer. For example, for GPU-equipped systems, CUDA-specialized PGs and OpenACC-specialized PGs are deployed.
@@
Figure 2 shows the directory structure used in the multi-agent configuration, illustrating how PM, SE, and PGs organize their workspaces.
@@ III-C2 Inter-agent Communication Mechanism
VibeCodeHPC utilizes a terminal multiplexer (terminal software that can manage multiple sessions) called Tmux to enable user monitoring of agent activities and communication between agents. Tmux provides functionality to simultaneously display multiple sessions divided into panes on a GUI and message sending functionality between sessions (tmux send-keys command). VibeCodeHPC utilizes “Tmux Multi-agent Communication,” [^13] https://github.com/nishimoto265/Claude-Code-Communication an inter-agent communication library for Claude Code that leverages this functionality. What communication each agent performs with other agents is described in the markdown configuration file that defines each agent’s operation, for example, ”PG must report execution results to PM”, ”When errors occur, report to SE and request guidance” and so on. These communication rules enable information sharing and collaborative operation between agents. Figure 3 shows an example of the desktop during multi-agent operation, where each agent runs in a separate tmux pane.
Figure 3: Desktop view during multi-agent operation with tmux panes
@@ III-C3 Agent Management and Dynamic Deployment by PM
When the system starts, the PM agent launches first and loads the markdown configuration file describing the PM’s role. The PM launches other agents according to its contents and manages subsequent operations. In existing multi-agent implementations (e.g., Tmux Multi-agent Communication Demo), a method was adopted that fixedly assigns agent configurations at startup. However, VibeCodeHPC implements a more flexible dynamic agent deployment functionality. The dynamic agent deployment functionality allows the PM to dynamically change the number and specialization of PGs according to situations such as:
@@ III-C4 Guaranteeing Persistent Agent Operation
LLM-based agents operate according to initially given instructions, but when the LLM’s context window (memory area) becomes full, problems arise where initial instructions are overlooked. In the case of Claude Code, when the context window becomes full, an automatic compression function called ”auto-compact” operates. This function compresses old information to secure memory capacity, but compression is not perfect. As a result, cases may occur where agents forget initial instructions or cause unexpected behavior leading to activity cessation. To address such situations, VibeCodeHPC implements the following mechanisms:
@@ IV-A Experimental Setup
We conducted experiments using the Supercomputer “Flow” Type II Subsystem [15] at the Information Technology Center, Nagoya University. The system specifications include: two Intel Xeon Gold 6230 CPUs (20 cores, 2.10–3.90 GHz); four NVIDIA Tesla V100 GPUs (2560 FP64 cores, up to 1530 MHz); 384 GiB DDR4 main memory (2933 MHz); and 32 GiB HBM2 device memory per GPU. The theoretical peak performance reaches 33.888 TFLOPS in double precision (1.344 TFLOPS per CPU socket, 7.8 TFLOPS per GPU). Memory bandwidth comprises 281.5 GB/s for main memory and 900 GB/s per GPU for device memory. GPU interconnection uses NVLINK2 (50 GB/s bidirectional), while CPU-GPU connection employs PCI-Express 3.0 (x16). GCC 4.8.5 and CUDA 11.2 are used for the compilation of generated codes. The following software is in use on the local side for VibeCodeHPC.
  • VibeCodeHPC-jp v0.7.5151515https://github.com/Katagiri-Hoshino-Lab/VibeCodeHPC-jp/releases/tag/v0.7.5
  • Windows 11 (local PC)
    • WSL version: 2.4.13.0
    • Kernel Version: 5.15.167.4-1
    • Ubuntu 22.04.5 LTS (Jammy Jellyfish)
    • tmux 3.2a
    • git version 2.34.1
    • gh version 2.4.0+dfsg1 (2022-03-23 Ubuntu 2.4.0+dfsg1-2)
  • OpenSSH_8.9p1 Ubuntu-3ubuntu0.13, OpenSSL 3.0.2 15 Mar 2022
  • nvm 0.39.7
  • Node.js v22.16.0
  • npm 10.9.2
  • Claude Code 2.0.8 – Claude Sonnet 4.5
  • Desktop Commander MCP v0.2.17
  • jq-1.6
  • Python 3.10.12
    • pip 22.0.2 from /usr/lib/python3/dist-packages/pip (python 3.10)
    • matplotlib 3.10.5
    • numpy 2.2.6
    • pandas 2.3.2
    • scipy 1.15.3
@@
The following common items were specified in the requirement definition document for both experiments. The optimization goal was set to maximize throughput, with the instruction to “continue optimization to the limit and aim for the highest performance on a single node.” The budget for computing resource consumption (utilization points based on job execution time and resource usage) was specified as: minimum consumption: 100 points, target: 500 points, upper limit: 1000 points. Figure 5 shows an example of budget tracking visualization from the matrix multiplication multi-agent EX3 experiment, automatically generated by parsing the ChangeLog.md maintained by agents. The following supplementary instructions were given to each agent: CD should frequently push all code except system information to GitHub; SE should exclude code that does not meet requirements when creating graphs; all agents should record both performance (GFLOPS) and execution time, with GFLOPS as the unified vertical axis. Note that the reported performance values include CPU-GPU data transfer time.

### v3

@@ I Introduction
In recent years, large language models (LLMs) have shown strong capabilities in automatic code generation, transforming natural language specifications into executable programs. Trained on large-scale source code corpora, these models can generate syntactically correct and semantically meaningful code and have been applied to software development tasks such as prototyping, bug fixing, refactoring, and testing. Prior work has demonstrated that large-scale code-trained models can generate effective solutions for programming problems [4], and that prompt engineering and in-context learning improve code correctness and robustness [2][26]. More recently, coding agent systems have emerged that autonomously perform user-specified tasks through iterative execution and refinement. Representative examples include Claude Code [^1] https://github.com/anthropics/claude-code, Codex CLI [^2] https://github.com/openai/codex, and Gemini CLI [^3] https://github.com/google-gemini/gemini-cli, which integrate LLMs with command-line interfaces (CLI). However, despite these advances, existing LLM-based systems primarily target general-purpose programming and do not sufficiently address performance-critical domains such as high-performance computing (HPC) applications on supercomputers, particularly with respect to performance optimization.
HPC programming imposes stringent requirements beyond functional correctness, including architecture-specific optimizations, parallelization across threads and processes, and trade-offs between numerical precision, performance, and power consumption. HPC applications typically rely on parallel programming models such as MPI, OpenMP, CUDA, and OpenACC, which must be selected according to application characteristics and execution environments. In addition, supercomputer usage requires system-specific operations, including batch job execution, software module management, and Secure Shell (SSH) access.
These complexities pose significant challenges for both human developers and code generation AI agents. LLMs, like humans, are constrained by limited context windows, and prolonged complex tasks may lead to forgotten instructions or degraded behavior. Multi-agent systems, in which multiple LLMs with specialized roles collaborate, offer a promising approach to mitigating these limitations by distributing cognitive load across agents. This advantage is particularly relevant for complex domains such as HPC programming. Moreover, defining rigorous requirements in advance is often difficult in such settings. To address this, the Vibe Coding paradigm has been proposed [19][28], in which programs are incrementally developed through interactive dialogue between humans and LLMs without strict upfront specifications. This paradigm is well suited to complex development domains, including HPC programming.
@@
Based on the above background, this study proposes VibeCodeHPC, a multi-agent system based on LLMs for the automatic tuning of HPC programs on supercomputers. VibeCodeHPC is built on top of Claude Code and offers an integrated framework for program development targeting supercomputing environments. Thanks to its sophisticated multi-agent mechanism, the system not only enables vibe coding as if working with a “specialized HPC programming team,” but also achieves fully automated tuning through autonomous operation that minimizes user intervention. To realize these objectives, VibeCodeHPC provides the following core functionalities:
@@
This paper is organized as follows. Section II reviews related work. Section III presents VibeCodeHPC and its implementation, focusing on multi-agent features. Section IV demonstrates automated code optimization through autonomous operation. Finally, Section V summarizes the findings and discusses future research directions.
@@ II Related Work
II-A HPC Code Generation and Optimization
@@
HPC code generation has traditionally relied on domain experts to manually implement performance-critical kernels using parallel programming models such as MPI, OpenMP, CUDA, and OpenACC. Compiler-assisted approaches [43][23] partially automate this process through loop transformations, memory layout optimization, and vectorization, while domain-specific languages (DSLs) [40][16] generate optimized kernels from high-level specifications. Although these methods can achieve high performance, they often require explicit user annotations or DSL expertise, limiting accessibility.
@@
Recent studies have begun evaluating LLMs for parallel programming models and HPC kernels, including OpenMP-centric models and MPI+OpenMP code generation in scalable workloads [3] [7]. Early evaluations of LLMs for OpenMP/MPI/CUDA/OpenACC kernels [10] [39] and LLM-assisted portability from Fortran to C/C++ [37] suggest that HPC-specific constraints substantially affect correctness and performance outcomes. These works provide concrete baselines against which VibeCodeHPC can be positioned.
@@
Practical case studies that combine LLM-based coding agents with GPU acceleration and testing pipelines offer direct evidence of feasibility in HPC contexts [14]. Additionally, HPC-oriented LLM systems and evaluations for software development provide a broader landscape for integration and operational support [46, 8, 11, 30].
II-B Auto-tuning and Iterative Refinement
@@
Auto-tuning (AT) has been widely adopted in HPC to explore parameter spaces and identify optimal configurations for target architectures. Frameworks such as ATLAS [42], OpenTuner [1], and GPTune [27] automate this process using search algorithms including Bayesian optimization, genetic algorithms, and multi-armed bandits. In addition, iterative refinement strategies, in which code is repeatedly modified based on performance feedback, have been studied in machine learning–guided compilation [5] [12]. While these approaches effectively optimize existing code, their integration with LLM-driven code synthesis remains limited. Comparative studies contrasting agentic LLM approaches with classical autotuning for loop transformations indicate that LLM-driven methods can complement or accelerate parameter search in certain settings [34]. These findings motivate the combination of structured tuning with multi-agent orchestration.
II-C Iterative Prompt in LLM-based Code Generation
@@
However, iteration can introduce vulnerabilities. Shukla et al. [35] reported a 37.6% increase in critical vulnerabilities after five iterations, highlighting the need for human oversight. From a structural perspective, Ren et al. [32] proposed Knowledge-driven Prompt Chaining (KPC) for exception handling, while Li et al. [25] introduced Structured CoT, improving Pass@1 accuracy by up to 13.79% on HumanEval. Ridnik et al. [33] demonstrated that test-driven, multi-stage iterative flows improved Pass@5 accuracy from 19% to 44%. Wei et al. [41] investigated CUDA code optimization using multi-agent systems without additional training. However, the role-sharing approach proposed in our study has not been evaluated for cross-language optimization. These studies constitute methodological investigations into iterative prompting. Research on AI-agent collaboration paradigms, as exemplified by VibeCodeHPC, remains nascent.
II-D Multi-Agent Frameworks and Orchestration Tools
Beyond research papers, the recent surge of general-purpose multi-agent frameworks (e.g., AutoGen[44], MetaGPT[13], CAMEL[24], and Magentic-One[9]) and agent orchestration tools (Claude Code subagents [^4] https://docs.anthropic.com/en/docs/claude-code/sub-agents, SuperClaude [^5] https://github.com/NomenAK/SuperClaude, Claude-Flow [^6] https://github.com/ruvnet/claude-flow, Claude Squad [^7] https://github.com/smtg-ai/claude-squad, Vibe Kanban [^8] https://github.com/BloopAI/vibe-kanban, KAMUI [^9] https://3rd.kamui.ai/en) highlights the growing interest in collaborative LLM workflows. Microsoft’s agent orchestration patterns categorize Magentic orchestration as a distinct mode, which aligns closely with VibeCodeHPC’s collaborative, role-separated workflow; however, a framework specialized for supercomputer code development remains unproposed[29].
@@ II-F Summary
Existing research in LLM-based code generation has achieved progress in producing functionally correct programs, but it rarely addresses HPC-specific performance optimization. On the other hand, the HPC community has developed powerful auto-tuning and compiler-based optimization techniques, yet these methods often lack the flexibility of LLM-based code synthesis. Bridging these domains – by integrating LLM-driven generation, HPC performance tuning, and iterative improvement—remains an open challenge, which our proposed VibeCodeHPC framework seeks to address.
Figure 1: Workflow of VibeCodeHPC.
@@ III Overview and Implementation of VibeCodeHPC
This section explains the system overview of VibeCodeHPC, the operational flow from the user’s perspective, and the implementation realizing it. In particular, we provide explanation of the multi-agent implementation, which is the core innovation of VibeCodeHPC.
@@ III-A System Components
VibeCodeHPC consists of the following main components.
@@
    • SSH connection management: Establishment and maintenance of remote access to supercomputers
@@
    This information is provided to support agents in appropriately performing system-specific operations.
@@ III-B Workflow
Figure 1 illustrates the workflow of VibeCodeHPC, which is described as follows.
@@
    System Startup: Although VibeCodeHPC can be operated anywhere, considering the usage pattern of supercomputers, it is assumed to run on the user’s local Linux terminal that accesses the supercomputer, from which it launches Claude Code.
@@
    Monitoring and Result Confirmation: Users can monitor each agent’s operation in real time, and VibeCodeHPC automatically generates work logs that allow users to track optimization progress. When the user’s objectives are achieved, a final report is generated, summarizing the optimization process, achieved performance, and applied techniques.
@@ III-C1 Agent Configuration
VibeCodeHPC enables collaborative operation of multiple agents specialized for different tasks. Their roles are defined in markdown-format configuration files loaded by Claude Code at startup and can be customized.
@@
  • Programmer (PG): An agent responsible for actual programming tasks. PGs are deployed for each specialized field, such as CPU optimization, GPU optimization (CUDA, OpenACC), distributed parallelization using MPI, and algorithm optimization. Deployment templates are prepared according to the system configuration of the target supercomputer; for example, GPU-equipped systems deploy CUDA- and OpenACC-specialized PGs.
@@
Figure 2 shows the directory structure of the multi-agent configuration, illustrating the organization of PM, SE, and PG workspaces.
@@ III-C2 Inter-agent Communication Mechanism
VibeCodeHPC utilizes a terminal multiplexer (terminal software that can manage multiple sessions) called Tmux to enable user monitoring of agent activities and communication between agents. Tmux provides functionality to simultaneously display multiple sessions divided into panes on a GUI and message sending functionality between sessions (tmux send-keys command). VibeCodeHPC utilizes “Tmux Multi-agent Communication,” [^12] https://github.com/nishimoto265/Claude-Code-Communication an inter-agent communication library for Claude Code that leverages this functionality. What communication each agent performs with other agents is described in the markdown configuration file that defines each agent’s operation, for example, “PG must report execution results to PM,” “When errors occur, report to SE and request guidance” and so on. These communication rules enable information sharing and collaborative operation between agents. Figure 3 shows an example of the desktop during multi-agent operation, where each agent runs in a separate tmux pane.
Figure 3: Desktop view of multi-agent operation with tmux panes
@@ III-C3 Agent Management and Dynamic Deployment by PM
At system startup, the PM agent is launched first and loads a markdown configuration file describing its role. Based on this configuration, the PM launches other agents and manages subsequent operations. Existing multi-agent implementations (e.g., Tmux Multi-agent Communication Demo) typically fix agent configurations at startup, whereas VibeCodeHPC supports flexible dynamic agent deployment. This functionality allows the PM to dynamically adjust the number and specialization of PGs according to situations such as:
@@ III-C4 Guaranteeing Persistent Agent Operation
LLM-based agents operate according to initially given instructions; however, when the LLM’s context window becomes full, these instructions may be overlooked. In Claude Code, an automatic compression mechanism called auto-compact compresses older information to free memory, but this process is imperfect and may lead to forgotten instructions or unexpected behavior, including activity cessation. To address these issues, VibeCodeHPC implements the following mechanisms:
@@ IV-A Experimental Setup
We conducted experiments using the Supercomputer “Flow” Type II Subsystem [15] at the Information Technology Center, Nagoya University. The system specifications include: two Intel Xeon Gold 6230 CPUs (20 cores, 2.10–3.90 GHz); four NVIDIA Tesla V100 GPUs (2560 FP64 cores, up to 1530 MHz); 384 GiB DDR4 main memory (2933 MHz); and 32 GiB HBM2 device memory per GPU. The theoretical peak performance reaches 33.888 TFLOPS in double precision (1.344 TFLOPS per CPU socket, 7.8 TFLOPS per GPU). Memory bandwidth comprises 281.5 GB/s for main memory and 900 GB/s per GPU for device memory. GPU interconnection uses NVLINK2 (50 GB/s bidirectional), while CPU-GPU connection employs PCI-Express 3.0 (x16). GCC 4.8.5 and CUDA 11.2 are used for the compilation of generated codes. The local environment for VibeCodeHPC consists of: VibeCodeHPC-jp v0.7.5141414https://github.com/Katagiri-Hoshino-Lab/VibeCodeHPC-jp/releases/tag/v0.7.5, Windows 11 with WSL 2.4.13.0 (Ubuntu 22.04.5 LTS, tmux 3.2a, git 2.34.1, gh 2.4.0), OpenSSH 8.9p1, Node.js v22.16.0, npm 10.9.2, Claude Code 2.0.8 (Claude Sonnet 4.5), Desktop Commander MCP v0.2.17, jq 1.6, and Python 3.10.12 (matplotlib 3.10.5, numpy 2.2.6, pandas 2.3.2, scipy 1.15.3).
@@
The following common items were specified in the requirement definition document for both experiments. The optimization goal was to maximize throughput, with instructions to continue optimization to the limit and achieve the highest possible single-node performance. The computing resource budget (utilization points based on job execution time and resource usage) was defined as 100 points minimum, 500 points target, and 1000 points upper limit. Figure 5 shows an example of budget tracking visualization from the matrix multiplication experiment (Multi EX3), automatically generated by parsing the ChangeLog.md maintained by agents. Supplementary instructions were also provided: CD should frequently push all non-system code to GitHub; SE should exclude non-compliant code when generating graphs; and all agents should record both performance and execution time, using GFLOPS as the unified vertical axis. Reported performance values include CPU–GPU data transfer time.
@@
Figure 6: Best performance progression for matrix multiplication across all experiments. Multi-agent runs (solid lines) generally achieve higher GFLOPS than solo-agent runs (dashed lines) within the 60-minute time limit.